OpenStudy (anonymous):

\[a_n =\frac{n!}{2^n}\]

OpenStudy (anonymous):

ratio test?

OpenStudy (anonymous):

\(\sum a_n\)?

OpenStudy (anonymous):

ratio test would work nicely, as you would have an avalanche of cancellation

OpenStudy (anonymous):

determine whether the sequence converges or diverges. if it converges, find the limit

OpenStudy (anonymous):

oh sequence, no series

OpenStudy (anonymous):

I'm going back to the first section of the chapter...trying to remember how I solved it

OpenStudy (anonymous):

\[lim_{n\rightarrow\infty}\]?

OpenStudy (anonymous):

is infinite

OpenStudy (anonymous):

let's see, how do I show this mathematically again?

OpenStudy (anonymous):

you could probably write down exactly what it is \[\frac{1\times 2\times 3\times 4\times ...\times n}{2\times 2\times 2\times 2\times ...\times 2}\]

OpenStudy (anonymous):

then maybe start at 3 and write \[\geq \frac{1}{2}\times \frac{3^{n-2}}{2^{n-2}}\] \[=\frac{1}{2}(\frac{3}{2})^{n-2}\] which pretty clearly goes to infinity since \(\frac{3}{2}>1\)

OpenStudy (anonymous):

oh perfect

OpenStudy (anonymous):

why n-2

OpenStudy (anonymous):

doesn't really matter because n goes to infinity, but i wrote \(n-2\) since i cut off the first two terms

OpenStudy (anonymous):

I see

OpenStudy (anonymous):

well thanks again Satellite!

OpenStudy (anonymous):

yw